Hi here is my question:
If I'm driving 12mph, how far did I drive in 2seconds?
Thanks!
ann

Answer by jim_thompson5910(35256)

60 seconds = 1 minute

60*60 seconds = 60*1 minute

3600 seconds = 60 minutes

3600 seconds = 1 hour

1 second = 1/3600 hours

2 seconds = 2/3600 hours

2 seconds = 1/1800 hours

-------------------------------------------------------

So if you drove for 2 seconds, then you also drove for 1/1800 hours


d = rt

d = 12*(1/1800)

d = 12/1800

d = 1/150

So you drove 1/150 miles, which converts to (1/150)*(5280 ft/1 mi) = 5280/150 = 35.2 feet

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------